# GQSA: Group Quantization and Sparsity for Accelerating Large Language Model Inference

### v3

@@ Abstract
Model compression has emerged as a mainstream solution to reduce memory usage and computational overhead. This paper presents Group Quantization and Sparse Acceleration (GQSA), a novel compression technique tailored for LLMs. Traditional methods typically focus exclusively on either quantization or sparsification, but relying on a single strategy often results in significant performance loss at high compression rates. In contrast, GQSA integrates quantization and sparsification in a tightly coupled manner, leveraging GPU-friendly structured group sparsity and quantization for efficient acceleration. Building upon system-algorithm co-design principles, we propose a two-stage sparse optimization strategy that ensures the performance superiority of the compressed model. On the engine side, we introduce a "task-centric" parallel strategy, which, to the best of our knowledge, is the first application in the domain of sparse computing. Compared to the traditional 2:4 sparse method, the GQSA offers a more flexible and adjustable sparsity rate, as well as a higher weight compression rate, and is efficiently compatible with weight-only quantization methods. Experimental results demonstrate that, under the GQSA W4S50% compression setting, the model’s accuracy surpasses that of both 2:4 pruning and W2 quantization. Furthermore, at the inference level, GQSA outperforms W2 by 1.26$\times$ and 2:4 pruning by 2.35$\times$ in terms of speed.
@@ 1 Introduction
To address these challenges, we propose a novel model compression method called GQSA, designed specifically for the decoding process and efficiently compatible with weight-only per-group quantization. GQSA explores a group sparsity pattern beyond the conventional 2:4 sparsity, achieving a better trade-off between accuracy and speed through a combination of algorithm-level optimizations and a customized software engine. Specifically, we reinterpret weight pruning as a particular form of quantization and introduce a group pruning based on group quantization. Our method incorporates the Block Sparse Row (BSR) format and designs a compact, low-precision weight storage structure to maximize the compression benefits of pruning and quantization. The GQSA method consists of two main stages. The first stage, Block Quantization-Pruning Optimization (BQPO), calibrates model parameters at the block level by optimizing weight distributions within block to minimize performance loss caused by group pruning and quantization. In the second stage, End-to-End Optimized Quantization-Pruning (E2E-OQP), the backbone network’s weights are frozen, and only the quantization parameters are fine-tuned to optimize the global network performance. Unlike BQPO, E2E-OQP considers the global error distribution across blocks. Freezing the backbone network can not only reduce memory usage but also improve training efficiency. Extensive experiments demonstrate that GQSA achieves significant advantages in both model accuracy and inference speed, especially when applied to newly released advanced models such as LLaMA-3 and LLaMA-3.1 model family.
@@
  • We propose a sparse scheme seamlessly compatible with widely used weight-only per-group quantization, effectively accelerating GEMV operations and reducing memory usage.
@@ 3.2 GQS Layer
where rowIndex[i] represent the offset of each row i, where i belongs to the range $[0,\text{rows}]$. The difference rowIndex[r+1] - rowIndex[i] indicates the number of non-zero groups in the i-th row. Additionally, rowIndex[rows] represents the total number of non-zero groups. The array groups[i] stores the indices of the non-zero groups; for instance, if groups[1] = 0, it means that the second group is located in the 0th column (in terms of group units). Finally, values stores the values of the non-zero groups for each row.
@@ 3.3 BQPO
In the first stage (shown in Figure 3(b)), we apply the BQPO method to optimize the GQS model, aiming to mitigate the accuracy loss caused by group quantization and pruning by adjusting the weight parameters within each block. Traditional QAT methods typically optimize the entire network’s weights in an end-to-end manner, as described by Equation (2) and (3). Similarly, most pruning methods adopt an end-to-end strategy to globally update unpruned parameters. However, these methods often require substantial computing resources and large-scale training datasets. For instance, pruning the LLaMA-65B model using LLM-Pruner and LoRAPruner requires 154GB and 72GB of memory, respectivelyZhang et al. (2024). To enhance optimization efficiency, BQPO employs a block-wise optimization strategy. Previous studies, such as OmniQuant and AffineQuant, have demonstrated that block-wise optimization can significantly reduce training time and memory usage. However, unlike OmniQuant and AffineQuant, which primarily focus on optimizing quantization parameters such as inter-channel smoothing factors and weight clipping thresholds, GQSA experiences more significant performance degradation due to its high sparsity in structured pruning and low-precision quantization. Therefore, BQPO focuses on optimizing the unpruned weights to restore model performance under extreme compression conditions. Due to the block-wise optimization strategy, this approach minimizes additional training costs, requiring only 26GB of memory for training.
@@ 3.4 E2E-OQP
Compared to BQPO, E2E-OQP not only emphasizes intra-block optimization but also considers the overall error across the entire network, thereby capturing cross-block information. As shown in Figure 3 (b), E2E-OQP differs from traditional QAT methods. We assume that BQPO has already produced a well-optimized model in the first stage; thus, E2E-OQP uses BQPO-optimized weights as initialization. During training, we freeze the main network weights $\tilde{W}$ and optimize only the quantization parameters $s$ and $z$ to further fine-tune model performance. The implementation of E2E-OQP highlights the advantages of the GQSA approach. When fine-tuning the quantization parameters, using the BSR format, we quantize the remaining group weights to low bits and freeze them, while the pruned groups are no longer retained. This strategy allows us to effectively fine-tune the quantization parameters and restore the performance of the GQSA model without introducing sparse masks. Overall, E2E-OQP significantly reduces memory usage by focusing exclusively on optimizing the quantization parameters of the remaining group and maintaining the main network at 4-bit quantization. This approach requires only 64GB of memory, which is substantially less than the memory demanded by LoRAPrune.
@@ 3.5 Custom Software Engine
Figure 5: Workload balancing through parallel task partitioning.
@@ 4 Experiments
In this section, we provide a comprehensive evaluation of GQSA, focusing on both model performance and inference speed, and demonstrate its advantages over existing model compression techniques through extensive experiments.
@@ 4.1 Experimental Settings
Models and Tasks. We selected the LLaMA Touvron et al. (2023a), LLaMA-2Touvron et al. (2023b), LLaMA-3, LLaMA-3.1Dubey et al. (2024) and OPTZhang et al. (2022) models to benchmark our method. Following previous studies, we evaluated the model’s language modeling capability on the WikiText2Merity et al. (2016) and C4Raffel et al. (2020) datasets. To assess performance on zero-shot tasks, we selected several mainstream benchmarks, including PIQABisk et al. (2020), ARCClark et al. (2018), HellaSwagZellers et al. (2019), and WinograndeSakaguchi et al. (2021), and conducted evaluations using lm-eval [^1] https://github.com/EleutherAI/lm-evaluation-harness.
@@ 4.2 Evaluation on Language Generation Tasks
To assess the performance of GQSA under extreme compression conditions, we first compared its perplexity against baseline method. As shown in Table 1, GQSA surpasses the performance of current state-of-the-art weight-only per-group quantization methods, including GPTQ, QUIP, OmniQuant, LeanQuant, under a 50% structured pruning combined with INT4 quantization. It also surpasses mixed-precision quantization models like PB-LLM and SliM-LLM. Furthermore, GQSA achieves comparable results to 2:4 semi-structured pruning while delivering substantial improvements in compression ratio and speedup. Similar results are presented in Table 9 for OPT models (ranging from 1.3B to 13B parameters), where GQSA consistently matches or surpasses baseline methods, even under more stringent compression settings. Furthermore, we observe that existing model compression methods often experience significant performance degradation on the latest large language models (e.g., LLaMA-3 and LLaMA-3.1). In contrast, GQSA demonstrates robust performance even in scenarios where other methods encounter substantial performance degradation.
@@ 5 Conclusion
Appendix A The effects of BQPO and E2E-OQP on the model’s performance
@@
Appendix B A comparison of the effects of pruning and quantization on inference performance
@@
Appendix C Combining the advantages of structured pruning and group quantization
@@
Appendix D Inference throughput of GQSA
@@
Appendix E Differences from Sparse Methods in Traditional CNNs
@@
Appendix F The advantages of group quantization compared to standard per-layer or per-label quantization methods.
@@
Appendix G Results of GQSA on the OPT model
To verify the generalization ability of GQSA on different model families, we conduct experiments on OPT models (ranging from 1.3B to 13B parameters). Table  9 shows similar results to the LLaMA model family, where GQSA consistently matches or outperforms the baseline methods even under stricter compression settings.
@@ Appendix G Results of GQSA on the OPT model
Appendix H GQSA inference latency and memory consumption

### v4

@@ Abstract
Model compression has emerged as a mainstream solution to reduce memory usage and computational overhead. This paper proposes GQSA, a novel model compression framework specifically designed for LLMs. Traditional methods typically focus exclusively on either quantization or sparsification, but relying on a single strategy often results in significant performance loss at high compression rates. In contrast, GQSA integrates quantization and sparsification in a tightly coupled manner, leveraging GPU-friendly structured group sparsity and quantization for efficient acceleration. Building upon system-algorithm co-design principles, we propose a two-stage sparse optimization strategy that ensures the performance superiority of the compressed model. On the engine side, we introduce a "task-centric" parallel strategy, which, to the best of our knowledge, is the first application in the domain of sparse computing. Compared to the traditional 2:4 sparse method, the GQSA offers a more flexible and adjustable sparsity rate, as well as a higher weight compression rate, and is efficiently compatible with weight-only quantization methods. Experimental results demonstrate that, under the GQSA W4S50% compression setting, the model’s accuracy surpasses that of both 2:4 pruning and W2 quantization. Furthermore, at the inference level, GQSA outperforms W2 by 1.26$\times$ and 2:4 pruning by 2.35$\times$ in terms of speed.
@@ 1 Introduction
To address these challenges, we propose a novel model compression method called GQSA, designed specifically for the decoding process and efficiently compatible with weight-only per-group quantization. GQSA explores a group sparsity pattern beyond the conventional 2:4 sparsity, achieving a better trade-off between accuracy and speed through a combination of algorithm-level optimizations and a customized software engine. Specifically, we reinterpret weight pruning as a particular form of quantization and introduce a group pruning based on group quantization. Our method incorporates the Block Sparse Row (BSR) format and designs a compact, low-precision weight storage structure to maximize the compression benefits of pruning and quantization. The GQSA method consists of two main stages. The first stage, Block Quantization-Pruning Optimization (BQPO), calibrates model parameters at the block level by optimizing weight distributions within block to minimize performance loss caused by group pruning and quantization. In the second stage, End-to-End Optimized Quantization-Pruning (E2E-OQP), the backbone network’s weights are frozen, and only the quantization parameters are fine-tuned to optimize the global network performance. Unlike BQPO, E2E-OQP considers the global error distribution across blocks. Freezing the backbone network can not only reduce memory usage but also improve training efficiency. Extensive experiments demonstrate that GQSA achieves significant advantages in both model accuracy and inference speed, especially when applied to newly released advanced models such as LLaMA-3 and LLaMA-3.1 model family and Qwen2.5 models.
@@
  • We propose a sparse scheme seamlessly compatible with widely used weight-only and weight-activation quantization, effectively accelerating GEMV operations and reducing memory usage.
@@ 3 GQSA
In this section, we provide a detailed exposition of GQSA. Section 3.1 begins with an examination of weight quantization and salient weight selection principles. Building upon these foundations, Section 3.2 introduces the innovative GQS Layer, designed to maximize the compression advantages from both quantization and pruning. The subsequent sections ( 3.3 and  3.4) detail our two-stage optimization algorithm, which delivers exceptional model performance. Concluding the section,  3.5 proposes a novel task-centric parallel strategy for efficient inference acceleration.
@@ 3.2 GQS Layer
[…]
@@ 3.3 BQPO
In the first stage (Figure 2(b)), we apply the BQPO method to optimize the GQS model, aiming to mitigate the accuracy degradation caused by group quantization and pruning. This is achieved by adjusting the weight parameters within each block. Traditional QAT methods typically optimize the entire network’s weights in an end-to-end fashion, as illustrated in Equations (2) and (3). Similarly, most pruning approaches adopt a global end-to-end strategy to update the remaining unpruned parameters. However, such methods often demand substantial computational resources and large-scale training datasets. To improve optimization efficiency, BQPO adopts a block-wise optimization strategy. Prior studies, such as OmniQuant and AffineQuant, have shown that block-wise optimization can significantly reduce both training time and memory consumption. Unlike OmniQuant and AffineQuant, which primarily optimize quantization parameters (inter-channel smoothing factors and weight clipping thresholds), GQSA suffers from more severe performance degradation due to its combination of high structured sparsity and low-precision quantization. As a result, BQPO focuses on optimizing the remaining weights to recover performance under extreme compression settings. This block-wise approach enables significant performance restoration with minimal additional training cost compared to global optimization techniques.
@@ 3.4 E2E-OQP
Compared to BQPO, E2E-OQP not only performs intra-block optimization but also accounts for the overall error across the entire network, thereby capturing cross-block dependencies. As illustrated in Figure 2(b), E2E-OQP differs from conventional quantization-aware training (QAT) methods. Assuming that BQPO has already yielded a well-optimized model in the first stage, E2E-OQP initializes training using the BQPO-optimized weights. During this phase, we freeze the primary network weights $\tilde{W}$ and optimize only the quantization parameters $s$ and $z$ to further refine model performance. The design of E2E-OQP underscores the advantages of the GQSA framework. Specifically, during fine-tuning, we employ the block-sparse row (BSR) format: the remaining group weights are quantized to low bit-width and frozen, while pruned groups are discarded entirely. This strategy enables effective fine-tuning of the quantization parameters without requiring sparse masks, thereby restoring the performance of the GQSA model under extreme compression. Overall, E2E-OQP achieves substantial memory savings by focusing solely on the quantization parameters of the remaining groups while maintaining 4-bit quantization across the main network. A detailed comparison of the resource consumption of BQPO and E2E-OQP is provided in Appendix A, demonstrating the efficiency advantages of the GQSA approach.
Figure 5: Workload balancing through parallel task partitioning.
@@ 4.1 Experimental Settings
Models and Tasks. We selected the LLaMA  Touvron et al. (2023a), LLaMA-2 Touvron et al. (2023b), LLaMA-3, LLaMA-3.1 Dubey et al. (2024) and OPT Zhang et al. (2022) models to benchmark our method. Following previous studies, we evaluated the model’s language modeling capability on the WikiText2 Merity et al. (2016) and C4 Raffel et al. (2020) datasets. To assess performance on zero-shot tasks, we selected several mainstream benchmarks, including PIQA Bisk et al. (2020), ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), and Winogrande Sakaguchi et al. (2021), and conducted evaluations using lm-eval.
@@ 4.2 Evaluation on Language Generation Tasks
To assess the performance of GQSA under extreme compression conditions, we first compared its perplexity against baseline method. As shown in Table 1, GQSA surpasses the performance of current state-of-the-art weight-only per-group quantization methods, including GPTQ, QuIP, OmniQuant, LeanQuant, under a 50% structured pruning combined with INT4 quantization. It also surpasses mixed-precision quantization models like PB-LLM and SliM-LLM. Furthermore, GQSA achieves comparable results to 2:4 semi-structured pruning while delivering substantial improvements in compression ratio and speedup. Similar results are presented in Table 14 for Qwen2.5 and Table 15 for OPT models, where GQSA consistently matches or surpasses baseline methods, even under more stringent compression settings. Furthermore, we observe that existing model compression methods often experience significant performance degradation on the latest large language models (e.g., LLaMA-3 and LLaMA-3.1). In contrast, GQSA demonstrates robust performance even in scenarios where other methods encounter substantial performance degradation.
@@ 5 Conclusion
Appendix A Training Efficiency of GQSA
Table 5 lists the memory and time required to train the Lllama-2 model using GQSA. The results show that GQSA requires only minimal resource overhead, and the 7B model only takes less than 10 hours to train with 9.3GB of memory, which is much less than the 14GB memory requirement to load the FP16 model, which is very efficient. It is also significantly better than other pruning methods, such as LLM-Pruner, which requires 18GB and takes less training time.
LLaMA-2 BQPO E2E-OQP
Memory (GB) Time (h) Memory (GB) Time (h)
7B 9.3 5.1 7.6 4.2
13B 14.3 7.3 11.7 6.4
Table 5: Detailed training time and training memory for GQSA at different model sizes and quantization bits on a single A100-40GB GPU.
Appendix B The effects of BQPO and E2E-OQP on the model’s performance
@@
Appendix C GQSA performance under weight-activation quantization
Unlike other algorithms that are limited to weight-only quantization, weight-activation quantization or model pruning, GQSA can not only efficiently combine pruning with weight-only quantization, but also support pruning with weight activation quantization. Our GPU-friendly grouped semi-structured sparse solution can be seamlessly combined with weight-only quantization or weight+activation quantization. On the basis of quantization, we can further improve performance by skipping some operations through sparsity.
Model Settings WikiText2 C4
LLaMA-2-7B W4A8S50% 7.84 11.04
LLaMA-2-13B W4A8S50% 14.09 21.26
Table 7: Performance comparison of GQSA with weight-activation quantization.
As show in Table 7, GQSA effectively preserves model accuracy under W4A8S50% quantization for both LLaMA-2-7B and LLaMA-2-13B architectures, maintaining strong performance despite simultaneous weight and activation quantization with 50% sparsity.
Appendix D Comparison of Quantized & Pruned Works
A comparison with SparseGPT’s joint sparsification and quantization. In SparseGPT’s report, "Joint Sparsification & Quantization" performs worse than "Sparsification-only," so we initially did not include it in our main content. However, for completeness, the following Table 8 presents a direct comparison between GQSA and SparseGPT’s "Joint Sparsification & Quantization" on LLaMA-2-13B and LLaMA-3-8B. The results demonstrate that GQSA provides a more significant performance advantage.
Method LlaMA-2-13B LLaMA-3-8B
WikiText2 C4 WikiText2 C4
SparseGPT 2:4 8.32 11.30 16.56 22.99
SparseGPT 2:4+INT4 9.25 12.74 19.43 26.34
GQSA W4S50% 7.80 10.93 13.84 20.85
Table 8: Performance comparison of GQSA with SparseGPT.
Comparison with contemporaneous works. To further validate the superiority of GQSA, we conduct comparative evaluations with contemporaneous methods including SliM-LoRA Mozaffari and Dehnavi (2024) and DC-W8A8 Wang et al. (2025). SliM-LoRA employs 4-bit weight quantization combined with Wanda’s 2:4 pruning but fails to overcome the limitations of semi-structured sparsity. Since NVIDIA’s 2:4 Tensor Cores do not support weight-only quantization, the inference acceleration benefits remain limited. Additionally, the 2:4 sparse format retains randomly positioned neurons and requires storing an equal amount of metadata to record their locations, preventing effective memory compression. SliM-LoRA also introduces the LoRA-Adapter; however, due to the quantization and sparsity of the main network, the LoRA-Adapter cannot be directly integrated and must be stored separately, increasing inference complexity. According to the SliM paper, its sparse quantization matrix can even reduce inference speed on the A100 GPU. In contrast, GQSA’s sparse quantization achieves a 4.3 $\times$ inference speedup over FP16, highlighting SliM’s shortcomings in both inference acceleration and memory compression. DC-W8A8 incorporates sparsity into W8A8 quantization but relies on unstructured sparsity with a sparsity rate of only 20%, offering minimal memory compression benefits. As stated in its paper, DC-W8A8 achieves only a 2.2$\times$ compression ratio compared to FP16, whereas GQSA achieves a 4.3$\times$ compression ratio. Moreover, GQSA significantly outperforms DC-W8A8 in inference acceleration.
Method OPT LLaMA-2
6.7B 13B 7B 13B
SliM-LoRA 47.08 47.96 54.26 57.85
DC-W8A8 48.55 - 60.89 -
GQSA W4S50% 53.26 56.39 59.36 64.96
Table 9: Performance comparison between GQSA and contemporaneous methods.
Table 9 presents the comparative evaluation of average accuracy on zero-shot tasks across different methods. The experimental results demonstrate that GQSA consistently outperforms both SliM and DC-W8A8 in terms of overall performance. Furthermore, GQSA achieves superior acceleration ratios and compression rates, while maintaining competitive accuracy. These advantages make GQSA particularly suitable for edge-side inference scenarios, where both computational efficiency and model compactness are critical.
Appendix E A comparison of the effects of pruning and quantization on inference performance
@@
Appendix F Combining the advantages of structured pruning and group quantization
@@
Appendix G Comparison of GQSA with Vector Quantization
Some of the latest low-bit quantization methods, such as AQLM Egiazarian et al. (2024) and QuIP# Tseng et al. (2024), employ vector quantization (VQ), which differs from uniform quantization techniques like GQSA. VQ constructs codebooks by learning the underlying data distribution, enabling better data preservation and potentially higher model performance. However, VQ methods rely on pre-trained codebooks (e.g., the E8P codebook used in QuIP# and the multi-codebook scheme in AQLM), which introduce considerable computational overhead during both training and inference. This makes them less practical for real-world deployment.
Method WikiText2 C4 Tokens Per Second
QuIP# W2 6.06 8.07 71.09
AQLM W2 5.60 7.47 68.1
GQSA W4S50% 7.80 10.93 228.95
Table 12: Comparison between GQSA and Vector Quantization
In contrast, GQSA combines uniform quantization with high sparsity to enable efficient inference acceleration in practical scenarios. As show in Table 12, while it may slightly underperform VQ-based methods like QuIP# and AQLM in terms of accuracy, it significantly outpaces them in inference speed—achieving up to 3.3$\times$ the speed of vector quantization methods under a small accuracy trade-off. No single method perfectly balances accuracy and computational efficiency; GQSA prioritizes inference speed, accepting a minor compromise in model accuracy to achieve substantial gains in performance.
Appendix H Inference throughput of GQSA
@@
Appendix I Differences from Sparse Methods in Traditional CNNs
@@
Appendix J The advantages of group quantization compared to standard per-layer or per-label quantization methods.
@@
Appendix K Results of GQSA on the Qwen model
To verify the generalization ability of GQSA on different model families, we conduct experiments on Qwen models (base and instruct model). Table  14 shows similar results to the LLaMA model family, where GQSA consistently matches or outperforms the baseline methods even under stricter compression settings.
Appendix L Results of GQSA on the OPT model
To verify the generalization ability of GQSA on different model families, we conduct experiments on OPT models (ranging from 1.3B to 13B parameters). Table  15 shows similar results to the LLaMA and Qwen model family, where GQSA consistently matches or outperforms the baseline methods even under stricter compression settings.
Qwen2.5-7B Qwen2.5-14B Qwen2.5-7B-Instruct Qwen2.5-14B-Instruct
Setting Method WikiText2 C4 WikiText2 C4 WikiText2 C4 WikiText2 C4
GPTQ 29.22 89.12 23.08 55.43 46.03 289.92 48.42 38.37
W2 OmniQuant 14.49 22.78 11.98 17.81 17.26 26.37 12.95 17.97
sparsegpt 11.25 17.17 10.13 15.39 11.92 17.85 10.95 16.24
2:4 wanda 14.78 22.84 11.74 18.24 15.80 23.83 12.06 18.73
[…]
Table 14: Wikitext2 and C4 perplexity (↓) for Qwen2.5 models, with a context length of 2048.
@@
Appendix M GQSA inference latency and memory consumption